# MITP-25-047 Exploring Nonperturbative Behaviour of Moments and Cumulants in Quantum Theories

### v1

@@ 2.1 Generating functionals for correlation functions
$\gamma_{n}(x_{1},\ldots,x_{n})=\braket{0}{\phi(x_{1})\ldots\phi(x_{n})}{0}\,.$ (2.1)
@@
The partition function measures the response of the field dynamics to an external local source $J(x)$, which enters the equation of motion as an inhomogeneous source term. In other words, it quantifies the probability of the vacuum state in the asymptotic past remaining in the vacuum state in the asymptotic future, $Z[J]=\braket{0}{0}_{J}$, under the influence of a driving force $J$.
@@
$G_{1}(x)=\frac{\gamma_{1}(x)}{Z_{0}}\,,\enspace G_{2}(x,y)=\frac{\gamma_{2}(x,y)}{Z_{0}}-\frac{\gamma_{1}(x)\gamma_{1}(y)}{Z_{0}^{2}}\,,$ (2.5)
@@
Here, $B_{n,k}$ denote the partial exponential Bell polynomials and the nonconnected correlation functions correspond to vacuum expectation values of the form $\gamma_{n}=\braket{0}{\phi^{n}(x)}{0}$. Due to Poincaré symmetry, the choice of $x$ is arbitrary.
@@
$\braket{0}{\phi(x)\phi(y)}{0}\sim\frac{1}{\left\lvert{x-y}\right\rvert^{2}}\,,$ (2.8)
@@ 2.2 Quantum fields as operator-valued distributions
In a scenario where all spacetime points of local field operator insertions coincide, correlation functions are ill-defined and their divergent structure needs to be carefully examined. For instance, this can be done in axiomatic formulation of QFT, where quantum fields are not local operators; rather, they are operator-valued distributions [34, 35]. These distributions are linear functionals that map test functions to operators that act on the Hilbert space of states,
@@ 3 Analogies with Ordinary Integrals
as $\lvert{J}\rvert\to\infty$. Thus, through elementary complex analysis, we find that the analytic function $Z(J)$ is of order $2p/(2p-1)$. This order is rational for any value of $p\geq 2$, i.e. for any nontrivial self-interaction. Consequently, by Picard’s theorem, $Z(J)$ must have infinitely many roots in the complex plane (see, e.g., [51]). This completes our proof that the cumulants, corresponding to the connected correlation functions, grow factorially with $n$. With this understanding derived from ordinary integrals, let us now proceed to examine higher-dimensional quantum theories.
@@ 4.1 An effective saddle point approximation
In practice, modern QFT calculations evaluate the path integral perturbatively, i.e. as a series expansion around the vacuum, although they often involve only a limited number of local field insertions. However, this approach may fail as $n$ increases, since the saddle point approximation is typically performed around the vacuum $\phi=0$. Similar to Eq. 3.1, as $n$ grows the support of the integrand shifts away from the perturbative vacuum at the origin and extends towards infinity along the real axis, significantly diminishing the accuracy of the naive approximation. This limitation of the perturbative approach can also be understood through Feynman diagrams. Typically, the number of Feynman diagrams contributing to the $n$-th order of the perturbative expansion grows factorially, $c_{n}\sim n!$, highlighting the asymptotic nature of perturbation theory. Consequently, for large $n$, the combination of $n$ and coupling, $g^{2}n$, distinguishes the perturbative regime, where $g^{2}n\ll 1$, from the nonperturbative regime of the theory, where $g^{2}n\gtrsim 1$. This implies that a naive summation of Feynman diagrams in the latter regime does not yield an accurate and meaningful result for large $g^{2}n$. To remedy this issue, we instead follow a semiclassical approach and express the moments as
@@
Thus, $\varphi_{0}$ is proportional to the Green’s function of the Helmholtz operator in $d$ dimensions, $\varphi_{0}\propto r^{-d/2+1}K_{d/2-1}(r)$. Similarly, the effective action can be decomposed into a power series in $\epsilon$ through Eq. 4.11. At leading order, it involves only the field $\varphi_{0}$, and is given by
@@
We remark that this result should be understood within the context of a double scaling limit where $g^{2}\to 0$ and $n\to\infty$, while keeping $\epsilon$ fixed. In fact, Eq. 4.14 is a large-$n$ representation of the familiar factorial growth of tree-level perturbation theory, through Stirling’s approximation. For example, the exponential form of $\gamma_{n}$ agrees with the overlap of an off-shell quantum state (created by an off-shell field acting on the vacuum) and an arbitrary $n$-particle state, $\braket{n}{\varphi}{0}$. More precisely, it matches the tree-level computation of the latter in the double scaling limit [4, 5, 6, 7, 8]. Physically, this is expected, as in both scenarios, any tree-level approximation is dominated by the number of Feynman diagrams, which grow factorially, thus yielding the above behaviour through Stirling’s approximation (modulo appropriate powers of the coupling).
So far, Eq. 4.11 dictates a perturbative expansion around $\epsilon=0$, which we expect to yield an accurate approximation in the perturbative regime of the quantum theory, where $\epsilon\ll 1$. In principle, it is viable to systematically compute higher-order terms of the classical field configuration, $\varphi_{i}$, along with the corresponding corrections to the effective action. This would extend the range of validity of the saddle point approximation. However, going beyond perturbation theory into the nonperturbative domain where $\epsilon\gtrsim 1$, requires the resummation of a substantial number of terms in this expansion. While exploring this domain is feasible in CFTs (see [29, 30, 31]), scenarios involving a mass term present additional complications. For example, although formal solutions for $\varphi_{1}$ exist, determining the corresponding correction to the effective action in an arbitrary number of dimensions proves to be challenging. One of the simplest nontrivial examples where an explicit solution can be obtained is QM, which serves as a one-dimensional QFT. In this context, the first correction to the effective action is given by
@@ 4.2 Resummation of the effective action in quantum mechanics
Addressing the nonperturbative behaviour of correlation functions at short distances requires a resummation of the perturbative expansion of the effective action to all-orders. We expect this approach to yield the exact results, as there are no instantons present on the real axis in the Borel plane that could spoil the resummation. In QM, this resummation of the effective action is feasible, because instead of relying on a path integral formulation, we can compute the moments directly from the ground-state wave function,
@@
[…]
We have confirmed this result for self-interactions up to $p=10$. It exemplifies the familiar factorial growth of observables associated with high multiplicities or large quantum numbers, highlighting the asymptotic nature of perturbation theory. Furthermore, we find that this leading contribution factorises from the full expression, as indicated in Eqs. 4.19 and 4.20. This suggests that the remaining loop contributions can be effectively resummed in the large-$n$ limit, similar to the case of the QM analogue of high-multiplicity form factors, $\braket{n}{x}{0}$ [56, 57, 58]. The leading quantum corrections to the tree-level result indicate an exponential form,
@@
$\begin{split}F_{c}^{p=2}=&-\frac{\epsilon^{2}}{32}+\frac{\epsilon^{3}}{192}-\frac{35}{24576}\epsilon^{4}+\ldots+\frac{1}{n}\left(-\frac{5}{16}\epsilon^{2}+\frac{45}{512}\epsilon^{3}+\ldots\right)\\ &+\frac{1}{n^{2}}\left(\frac{371}{768}\epsilon^{2}-\frac{3913}{12288}\epsilon^{3}+\ldots\right)+\mathcal{O}\left(\frac{1}{n^{3}}\right)\,,\end{split}$ (4.27)
@@
It is remarkable that this exponentiation is exact to all orders in perturbation theory, which is not guaranteed (for a detailed discussion of exponential series representations see [56]).
@@
This expression precisely matches the leading-order correction to the effective action $S_{\mathrm{eff}}$ in Eq. 4.15, which we have obtained from an effective saddle point approximation. This not only validates our approach to some extent but also facilitates an immediate identification of the exponent function $F_{0}$ with the effective saddle point. In this context, it would be interesting to investigate whether the $1/n$-corrections of $F$ correspond to the quantum fluctuations around the effective saddle point. At the same time, our result demonstrates that each coefficient of $S_{\mathrm{eff}}$ (or equivalently $F_{0}$) effectively resums an infinite number of loops at large $n$. This also means that the closed form in Eq. 4.31 enables us to resum the formal series expansion, and obtain the analytic continuation of $F_{0}(\epsilon)$ which is valid everywhere in the complex $\epsilon$-plane. We omit its explicit form here, as it is lengthy and convoluted. Instead, we illustrate its shape in Fig. 3.
Clearly, $F$ becomes positive beyond a certain critical value of $\epsilon$, indicating that the moments $\gamma_{n}$ grow exponentially at high energies. More importantly, the growth rate appears to depend on the self-interaction power, with larger values of $p$ leading to a suppression of this growth. This behaviour is consistent with our findings for ordinary integrals (see Fig. 1). Focussing on the nonperturbative domain, where $\epsilon\gtrsim 1$, this observation is further supported by an asymptotic expansion as $\epsilon\to\infty$,
@@ 4.3 Asymptotic behaviour of cumulants
While we can study the cumulants’ behaviour for any value of $\epsilon$ in this manner, the evaluation of Eq. 2.7 becomes numerically unstable for large values of $\epsilon$, due to the combinatorial complexities arising from the addition and subtraction of large numbers. To address this instability, we numerically determine the ground state wave function with high precision and then directly integrate the correlation functions $\gamma_{n}$, for arbitrary values of $n$ and $g^{2}$. These agree with our perturbative results and yield robust results for the cumulants $G_{n}$ determined from Eq. 2.7. An example of both $\gamma_{n}$ and $G_{n}$ is illustrated in Fig. 4. Notably, the moments $\gamma_{n}$, shown in the left panel of Fig. 4, exhibit a strong dependence on the self-interaction term of the theory as $n\to\infty$. This behaviour aligns with our expectations in the nonperturbative regime, where we have established a suppression of the exponent $F$ through the self-interaction power $p$, given in Eq. 4.33. In contrast, the cumulants $G_{n}$ are largely independent of the self-interaction term, meaning they do not significantly depend on $p$. More precisely, as $n\to\infty$, they grow asymptotically as
@@
Moreover, our results are intriguing in the context of effective field theories (EFTs), which accommodate operators of arbitrary power beyond renormalisable QFT models. In this framework, our results suggest that the physical information of the quantum theory carried by the fully connected functions is only marginally dependent on the precise operator structure in the UV. That is, in the large-$n$ limit, all operators of the EFT contribute equally. In other words, in the nonperturbative regime of a quantum theory, we would not be able to detect the presence of higher-order interactions through the form factors discussed in this work. However, higher-dimensional QFTs introduce additional complications that must be carefully addressed. Notably, all coupling constants beyond a $\phi^{4}$ term are dimensionful in four dimensions, which introduces an additional energy scale above which higher-dimensional operators become effective. This must be accounted for in the double scaling limit. Furthermore, it is important to recognise that the moments and cumulants discussed in this work do not represent physical observables in a strict sense. In a more realistic setting, physical correlation functions would be measured through scattering processes in particle collisions. This scenario would involve field insertions at different spacetime points, introducing nontrivial kinematics that can lead to suppression of interaction rates through phase space factors at large $n$, as the quantum theory would otherwise violate unitarity.
@@ Appendix A Perturbative Corrections to the Effective Saddle Point
Let us now outline how to obtain corrections to the effective action beyond leading oder. Following our perturbative approach, we find the equations of motion for the first correction to the saddle point field configuration, $\varphi_{1}$, to be
@@
Here, all contributions on the right hand side are known, in principle. Since the self-interactions of the field appear only as inhomogeneous terms in the equations of motion, we can express a formal solution for $\varphi_{1}$ using the free Green’s functions $D$ in position space,
@@
In the case of QM, we have $d=1$, simplifying all multi-dimensional integrals to one-dimensional integrals. Here, the leading-order contribution to the saddle point configuration is given by

### v2

@@ 2.1 Generating functionals for correlation functions
$\gamma_{n}(x_{1},\ldots,x_{n})=\braket{0|\phi(x_{1})\ldots\phi(x_{n})|0}\,.$ (2.1)
@@
The partition function measures the response of the field dynamics to an external local source $J(x)$, which enters the equation of motion as an inhomogeneous source term. In other words, it quantifies the probability of the vacuum state in the asymptotic past remaining in the vacuum state in the asymptotic future, $Z[J]=\braket{0|0}_{J}$, under the influence of a driving force $J$.
@@
$G_{1}(x)=\frac{\gamma_{1}(x)}{Z_{0}}\,,\kern 5.0ptG_{2}(x,y)=\frac{\gamma_{2}(x,y)}{Z_{0}}-\frac{\gamma_{1}(x)\gamma_{1}(y)}{Z_{0}^{2}}\,$, (2.5)
@@
Here, $B_{n,k}$ denote the partial exponential Bell polynomials and the nonconnected correlation functions correspond to vacuum expectation values of the form $\gamma_{n}=\braket{0|\phi^{n}(x)|0}$. Due to Poincaré symmetry, the choice of $x$ is arbitrary.
@@
$\braket{0|\phi(x)\phi(y)|0}\sim\frac{1}{\left\lvert{x-y}\right\rvert^{2}}\,$, (2.8)
@@ 2.2 Quantum fields as operator-valued distributions
In a scenario where all spacetime points of local field operator insertions coincide, correlation functions are ill defined and their divergent structure needs to be carefully examined. For instance, this can be done in axiomatic formulation of QFT, where quantum fields are not local operators; rather, they are operator-valued distributions [34, 35]. These distributions are linear functionals that map test functions to operators that act on the Hilbert space of states,
@@ 3 Analogies with Ordinary Integrals
as $\lvert{J}\rvert\to\infty$. Thus, through elementary complex analysis, we find that the analytic function $Z(J)$ is of order 2p/(2p1)2p/(2p-1). This order is rational for any value of $p\geq 2$, i.e. for any nontrivial self-interaction. Consequently, by Picard’s theorem, $Z(J)$ must have infinitely many roots in the complex plane (see, e.g., [51]). This completes our proof that the cumulants, corresponding to the connected correlation functions, grow factorially with $n$. With this understanding derived from ordinary integrals, let us now proceed to examine higher-dimensional quantum theories.
@@ 4.1 An effective saddle point approximation
In practice, modern QFT calculations evaluate the path integral perturbatively, i.e. as a series expansion around the vacuum, although they often involve only a limited number of local field insertions. However, this approach may fail as $n$ increases, since the saddle point approximation is typically performed around the vacuum $\phi=0$. Similar to Eq. 3.1, as $n$ grows the support of the integrand shifts away from the perturbative vacuum at the origin and extends towards infinity along the real axis, significantly diminishing the accuracy of the naive approximation. This limitation of the perturbative approach can also be understood through Feynman diagrams. Typically, the number of Feynman diagrams contributing to the $n$-th order of the perturbative expansion grows factorially, $c_{n}\sim n!$, highlighting the asymptotic nature of perturbation theory. Consequently, for large $n$, the combination of $n$ and coupling, $g^{2}n$, distinguishes the perturbative regime, where $g^{2}n\ll 1$, from the nonperturbative regime of the theory, where $g^{2}n\gtrsim 1$. This implies that a naive summation of Feynman diagrams in the latter regime does not yield an accurate and meaningful result for large $g^{2}n$. To remedy this issue, we instead follow a semiclassical approach and express the moments as (see also [29, 30])
@@
Thus, $\varphi_{0}$ is proportional to the Green’s function of the Helmholtz operator in $d$ dimensions, $\varphi_{0}\propto r^{-d/2+1}K_{d/2-1}(r)$. Similarly, the effective action can be decomposed into a power series in $\epsilon$ through Eq. 4.11. At leading order, it involves only the field $\varphi_{0}$ and is given by
@@
We remark that this result should be understood within the context of a double scaling limit where $g^{2}\to 0$ and $n\to\infty$, while keeping $\epsilon$ fixed. In fact, Eq. 4.14 is a large-$n$ representation of the familiar factorial growth of tree-level perturbation theory, through Stirling’s approximation. For example, the exponential form of $\gamma_{n}$ agrees with the overlap of an off-shell quantum state (created by an off-shell field acting on the vacuum) and an arbitrary $n$-particle state, $\braket{n|\varphi|0}$. More precisely, it matches the tree-level computation of the latter in the double scaling limit [4, 5, 6, 7, 8]. Physically, this is expected, as in both scenarios, any tree-level approximation is dominated by the number of Feynman diagrams, which grow factorially, thus yielding the above behaviour through Stirling’s approximation (modulo appropriate powers of the coupling).
So far, Eq. 4.11 dictates a perturbative expansion around $\epsilon=0$, which we expect to yield an accurate approximation in the perturbative regime of the quantum theory, where $\epsilon\ll 1$. In principle, it is viable to systematically compute higher-order terms of the classical field configuration, $\varphi_{i}$, along with the corresponding corrections to the effective action. This would extend the range of validity of the saddle point approximation. However, going beyond perturbation theory into the nonperturbative domain, where $\epsilon\gtrsim 1$, requires the resummation of a substantial number of terms in this expansion. While exploring this domain is feasible in CFTs (see [29, 30, 31]), scenarios involving a mass term present additional complications. For example, although formal solutions for $\varphi_{1}$ exist, determining the corresponding correction to the effective action in an arbitrary number of dimensions proves to be challenging. One of the simplest nontrivial examples where an explicit solution can be obtained is QM, which serves as a one-dimensional QFT. In this context, the first correction to the effective action is given by
@@ 4.2 Resummation of the effective action in quantum mechanics
Addressing the nonperturbative behaviour of correlation functions at short distances requires a resummation of the perturbative expansion of the effective action to all orders. We expect this approach to yield the exact results, as there are no instantons present on the real axis in the Borel plane that could spoil the resummation. In QM, this resummation of the effective action is feasible because instead of relying on a path integral formulation, we can compute the moments directly from the ground-state wave function,
@@
$\gamma_{n}^{\mathrm{tree}}=\frac{\Gamma\left( \frac{n+1}{2}\right)}{\sqrt{\pi}}\,.$ (4.21)
We have confirmed this result for self-interactions up to $p=10$. It exemplifies the familiar factorial growth of observables associated with high multiplicities or large quantum numbers, highlighting the asymptotic nature of perturbation theory. Furthermore, we find that this leading contribution factorises from the full expression, as indicated in Eqs. 4.19 and 4.20. This suggests that the remaining loop contributions can be effectively resummed in the large-$n$ limit, similar to the case of the QM analogue of high-multiplicity form factors, $\braket{n|x|0}$ [56, 57, 58]. The leading quantum corrections to the tree-level result indicate an exponential form,
@@
[…]
@@
It is remarkable that this exponentiation is exact to all orders in perturbation theory, which is not guaranteed (for a detailed discussion of exponential series representations, see [56]).
@@
This expression precisely matches the leading-order correction to the effective action $S_{\mathrm{eff}}$ in Eq. 4.15, which we have obtained from an effective saddle point approximation. This not only validates our approach to some extent but also facilitates an immediate identification of the exponent function $F_{0}$ with the effective saddle point. In this context, it would be interesting to investigate whether the $1/n$-corrections of $F$ correspond to the quantum fluctuations around the effective saddle point. At the same time, our result demonstrates that each coefficient of $S_{\mathrm{eff}}$ (or equivalently $F_{0}$) effectively resums an infinite number of loops at large $n$. This also means that the closed form in Eq. 4.31 enables us to resum the formal series expansion and obtain the analytic continuation of $F_{0}(\epsilon)$, which is valid everywhere in the complex $\epsilon$-plane. We omit its explicit form here, as it is lengthy and convoluted. Instead, we illustrate its shape in Fig. 3.
Clearly, $F$ exhibits a turning point before becoming positive beyond a certain critical value $\epsilon_{0}$, where $F(\epsilon_{0})=0$. The turning point is reminiscent of the point of optimal truncation, $\epsilon_{\mathrm{opt}}$. It marks the order beyond which the individual terms of the asymptotic series expansion of $F$ start to grow rapidly, and a careful resummation becomes necessary. We find that its location only mildly depends on the self-interaction power, and it appears to asymptotically approach $\epsilon_{\mathrm{opt}}\simeq 2$ as $p$ grows. We have verified this behaviour for self-interactions up to $p=8$. Similarly, the root of $F$ marks the point $\epsilon_{0}$ where exponential growth for the moments $\gamma_{n}$ sets in. That is, if one chooses a weak coupling $g^{2}$, $\epsilon_{0}$ indicates the value of $n$ for which the moments start to grow exponentially, $n\gtrsim\epsilon_{0}/g^{2}$. While, to the best of our knowledge, an analytic expression for $\epsilon_{0}$ does not exist for any value of $p$, we observe that it is monotonically shifted towards larger values as $p$ grows. Again, we have explicitly verified this for self-interactions up to $p=8$. We expect that it asymptotically behaves as $\epsilon_{0}\to\infty$ as $p\to\infty$, as this scenario naively corresponds to the infinite square well potential. In this case, one can show explicitly that the moments are exponentially suppressed for large $n$, such that $F$ is necessarily negative everywhere. We conclude that the growth rate depends on the self-interaction power, with larger values of $p$ leading to a suppression of this growth. This behaviour is consistent with our findings for ordinary integrals (see Fig. 1). Focussing on the nonperturbative domain, where $\epsilon\gtrsim 1$, this observation is further supported by an asymptotic expansion as $\epsilon\to\infty$,
@@ 4.3 Asymptotic behaviour of cumulants
While we can study the cumulants’ behaviour for any value of $\epsilon$ in this manner, the evaluation of Eq. 2.7 becomes numerically unstable for large values of $\epsilon$ due to the combinatorial complexities arising from the addition and subtraction of large numbers. To address this instability, we numerically determine the ground state wave function with high precision and then directly integrate the correlation functions $\gamma_{n}$, for arbitrary values of $n$ and $g^{2}$. These agree with our perturbative results and yield robust results for the cumulants $G_{n}$ determined from Eq. 2.7. An example of both $\gamma_{n}$ and $G_{n}$ is illustrated in Fig. 4. Notably, the moments $\gamma_{n}$, shown in the left panel of Fig. 4, exhibit a strong dependence on the self-interaction term of the theory as $n\to\infty$. This behaviour aligns with our expectations in the nonperturbative regime, where we have established a suppression of the exponent $F$ through the self-interaction power $p$, given in Eq. 4.33. In contrast, the cumulants $G_{n}$ are largely independent of the self-interaction term, meaning they do not significantly depend on $p$. More precisely, as $n\to\infty$, they grow asymptotically as
@@
That said, it would be interesting to investigate the massless regime of our results, naively resembling the CFT limit of the quantum theory. In practice, this may be achieved through the framework of “exact perturbation theory,” where the perturbative ansatz for a theory with massive degrees of freedom can be suitably reorganised to obtain nonperturbative results for its massless limit [60, 61]. While this method, in principle, facilitates the precise computation of vacuum expectation values in the massless limit, their physical interpretation is somewhat more involved. On a technical level, the ultra-short distance limit of moments in a quantum theory featuring conformal symmetry may not be meaningful, as any test function providing a localised support for the field necessarily introduces a characteristic length scale, beyond which individual spacetime points become indistinguishable. This length scale breaks the conformal invariance of the quantum theory. More generally and strictly speaking, a CFT also does not feature the notion of a finite separation between quanta, such that there are no asymptotic in- and out-states to describe scattering processes. Consequently, connected correlation functions and their associated $S$-matrix elements cannot be properly defined.
On the other hand, the semiclassical expansion around the effective saddle point, as outlined above, is able to shed light on quantum theories featuring global symmetries with an associated Noether charge. Naively, this Noether charge corresponds to the number $n$ of operator insertions at the same spacetime point. In this case, the method may facilitate an interpolation between Feynman-diagrammatic calculations, valid for $\epsilon\ll 1$, and the superfluid regime for CFTs at large charge, $\epsilon\gg 1$ [29, 30]. Closely following this approach, it would be interesting to supplement these examples with a massive degree of freedom. In this case, even determining only the first few terms of the effective action may provide valuable insights into the dynamics of large-charge sectors of the quantum theory.
Finally, our results are intriguing in the context of effective field theories (EFTs), which accommodate operators of arbitrary power beyond renormalisable QFT models. In this framework, our results suggest that the physical information of the quantum theory carried by the fully connected functions is only marginally dependent on the precise operator structure in the UV. That is, in the large-$n$ limit, all operators of the EFT contribute equally. In other words, in the nonperturbative regime of a quantum theory, we would not be able to detect the presence of higher-order interactions through the form factors discussed in this work. However, higher-dimensional QFTs introduce additional complications that must be carefully addressed. Notably, all coupling constants beyond a $\phi^{4}$ term are dimensionful in four dimensions, which introduces an additional energy scale above which higher-dimensional operators become effective. This must be accounted for in the double scaling limit. Furthermore, it is important to recognise that the moments and cumulants discussed in this work do not represent physical observables in a strict sense. In a more realistic setting, physical correlation functions would be measured through scattering processes in particle collisions. This scenario would involve field insertions at different spacetime points, introducing nontrivial kinematics that can lead to suppression of interaction rates through phase space factors at large $n$, as the quantum theory would otherwise violate unitarity.
@@ Appendix A Perturbative Corrections to the Effective Saddle Point
Let us now outline how to obtain corrections to the effective action beyond leading order. Following our perturbative approach, we find the equations of motion for the first correction to the saddle point field configuration, $\varphi_{1}$, to be
@@
Here, all contributions on the right-hand side are known, in principle. Since the self-interactions of the field appear only as inhomogeneous terms in the equations of motion, we can express a formal solution for $\varphi_{1}$ using the free Green’s functions $D$ in position space,
@@
In the case of QM, we have $d=1$, simplifying all multidimensional integrals to one-dimensional integrals. Here, the leading-order contribution to the saddle point configuration is given by